OpenStudy (anonymous):

explain the distance formula. then use it to calculate the distance between a(1 1) and b(7 -7)

OpenStudy (dinamix):

ab= \[\sqrt{(7-1)^2+((-7)-1)^2}\]

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

need some help?

OpenStudy (anonymous):

well tagg me if you decide you want my help...

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

sure.. let me explain.. so the distance formula is used to find the distance between two points

OpenStudy (anonymous):

kinda self explanatory haha

OpenStudy (anonymous):

but then what is the distance between

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (dinamix):

\[\sqrt{36+64} = \sqrt{100}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

So the distance formula is just a^2+b^2=c^2.

OpenStudy (faiqraees):

@paaannndddaaa7 If you still need contact me i will explain in detail